OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1).

OpenStudy (anonymous):

USe point slope form formula: y-y1=m(x-x1) where m is slope and (x1,y1) is a point that passes through the graph

OpenStudy (zale101):

m is the slope so m=-2 x1=5 y1=-1 plug those into the point slope form y-y1=m(x-x1)

OpenStudy (zale101):

can u do that?

OpenStudy (anonymous):

yeah thanks guys